# SHViT: Single-Head Vision Transformer with Memory Efficient Macro Design

### v1

@@ 1 Introduction
Based on these findings, we introduce a Single-Head Vision Transformer (SHViT) based on memory-efficient design principles, as a new family of networks that run highly fast on diverse devices. Experiments demonstrate that our SHViT achieves state-of-the-art performance for classification, detection, and segmentation tasks in terms of both speed and accuracy, as shown in Fig. 1. For instance, our SHViT-S4 achieves 79.4% top-1 accuracy on ImageNet with throughput of 14283 images/s on an Nvidia A100 GPU and 509 images/s on an Intel Xeon Gold 5218R CPU @ 2.10GHz, outperforming EfficientNet-B0 [24] by 2.3% in accuracy, 69.4% in GPU inference speed, and 90.6% in CPU speed. Also, SHViT-S4 has 1.3% better accuracy than MobileViTv21.0 [21] and is 2.4 faster on iPhone12 mobile device. For object detection and instance segmentation on MS COCO [25] using Mask-RCNN [26] detector, our model significantly outperforms EfficientViT-M4 [27] by 6.2 AP and 4.9 AP with a smaller backbone latency on various devices.
@@ 2 Analysis and Method
| Macro design |
Reso. |
@@ 2.1 Analysis of Redundancy in Macro Design
To substantiate our hypothesis, inspired by [10, 32, 27], we adopt 1616 patchify stem and 3-stage design. We build two models based on the MetaFormer block [28] and the two aforementioned macro designs (see Fig. 2 for details). Specifically, we configured both models to have a similar number of channels for equivalent feature map size. Surprisingly, model (b) is 3.0 / 2.8 faster on the GPU / CPU, respectively, although it performs 1.5% worse than (a). Furthermore, when trained at a resolution of 256256, (b) is not only comparable to (a) but also significantly faster.
@@
1) token representations with large receptive fields and reduced spatial redundancy can be utilized at the early stages. 2) It can diminish the feature map size by up to 16 times, leading to a significant reduction in memory access costs. 3) due to the aggressive stride design, there’s only a mild decrease in throughput when the resolution is increased, leading to effective performance enhancement (as shown in Fig. 2(b), 8, Tab. 2).
@@ 2.2 Analysis of Redundancy in Micro Design
First of all, in the early stages (Fig. 3 (a)), we observe that the top-performing heads tend to operate in a convolution-like manner, while heads that have minimal impact on performance when removed typically process information more globally. Also, as shown in Fig. 2 (b), the model using attention layers in the first stage exhibits a less favorable speed-accuracy trade-off compared to those employing depthwise convolution layers in the first stage. Hence, for efficiency, we use convolutions with spatial inductive bias as the token mixer in the initial stage.
@@ 2.3 Single-Head Self-Attention
| (a) : Bottleneck |
10.6 |
802 |
27195 |
73.6 |
| (b) : Full channel |
10.4 |
724 |
26432 |
75.1 |
| (c) : Partial channel (Ours) |
11.4 |
951 | 26878 | 75.2 |
@@ 3.1 Implementation Details
We conduct image classification on ImageNet-1K [23], which includes 1.28M training and 50K validation images for 1000 categories. All models are trained from scratch using AdamW [57] optimizer for 300 epochs with a learning rate of 10 and a total batch size of 2048. We use cosine learning rate scheduler [58] with linear warmup for 5 epochs. Weight decays are set to 0.025/0.032/0.035/0.03 for SHViT-S1 to S4. For fair comparison, we follow the same data augmentation proposed in [31], including Mixup [59], random erasing [60], and auto-augmentation [61]. For 384 and 512 resolution, we finetune the model for 30 epochs with weight decay of 10 and learning rate of 0.004. Additionally, we assess throughput performance across various hardware platforms. We measure GPU throughput on an Nvidia A100 with batch size of 256. For CPU and , we evaluate the runtime on an Intel(R) Xeon(R) Gold 5218R CPU @ 2.10GHz processor , with batch size of 16 (using a single thread). For , we convert the models to ONNX [62] runtime format. Mobile latency is measured using iPhone 12 with iOS version 16.5. We export the models (batch size is set to 1) using CoreML tools [63] and report the median latency over 1,000 runs. We also validate our model as an efficient vision backbone for object detection and instance segmentation on COCO [25] with RetinaNet [64] and Mask R-CNN [26], respectively. All models are trained under 1 schedule (12 epochs) following [40] on mmdetection library [65].
@@ 3.3 SHViT on downstream tasks
As shown in the above results, the large-stride patchify stem with 3-stage reduces not only computational costs but also generates meaningful token representations, especially at higher resolutions. Furthermore, the marked performance gap with EfficientViT, using a similar macro design, proves the efficacy of our micro design choices.
@@ 4 Related Work
Also, recent works [34, 35, 36, 37, 38, 39, 75, 27, 76] have demonstrated that numerous heads function in similar ways and can be pruned without notably affecting performance. EfficientViT [27] proposes feeding attention heads with different splits of the full channel to improve attention diversity. In addition, [76] presents a regularization loss for attention similarity. As opposed to reducing multi-head redundancy, we design module with single-head configuration, which not only inherently prevents multi-head redundancy but also saves computation costs.
@@ 5 Conclusion
In this work, we have investigated redundancies at both the spatial and channel dimensions of the architectural design commonly used by many established models. We then proposed 1616 patch embeddings with 3-scale hierarchical representations and Single-Head Self-Attention to address the computational redundancies. We further present our versatile Single-Head Vision Transformer, built upon our proposed macro/micro designs, that achieves ultra-fast inference speed and high performance on diverse devices and vision tasks.
@@
While our macro design is effective, there is a need for fine-grained (high-resolution) features to enhance performance further or to recognize small objects. Therefore, our future work focuses on devising cost-effective ways to utilize them. Integrating the single-head design into existing sophisticated attention mechanisms will be an intriguing way to explore.

### v2

@@ 1 Introduction
Based on these findings, we introduce a Single-Head Vision Transformer (SHViT) based on memory-efficient design principles, as a new family of networks that run highly fast on diverse devices. Experiments demonstrate that our SHViT achieves state-of-the-art performance for classification, detection, and segmentation tasks in terms of both speed and accuracy, as shown in Fig. 1. For instance, our SHViT-S4 achieves 79.4% top-1 accuracy on ImageNet with throughput of 14283 images/s on an Nvidia A100 GPU and 509 images/s on an Intel Xeon Gold 5218R CPU @ 2.10GHz, outperforming EfficientNet-B0 [24] by 2.3% in accuracy, 69.4% in GPU inference speed, and 90.6% in CPU speed. Also, SHViT-S4 has 1.3% better accuracy than MobileViTv21.0 [21] and is 2.4 faster on iPhone12 mobile device. For object detection and instance segmentation on MS COCO [25] using Mask-RCNN [26] detector, our model significantly outperforms EfficientViT-M4 [27] by 6.2 APbox and 4.9 APmask with a smaller backbone latency on various devices.
@@ 2 Analysis and Method
| Macro design | Reso. |
@@ 2.1 Analysis of Redundancy in Macro Design
To substantiate our hypothesis, inspired by [10, 32, 27], we adopt 1616 patchify stem and 3-stage design. We build two models based on the MetaFormer block [28] and the two aforementioned macro designs (see Fig. 2 for details). Specifically, we configured both models to have a similar number of channels for equivalent feature map size. Surprisingly, model (b) is 3.0 / 2.8 faster on the GPU / CPU, respectively, although it performs 1.5% worse than (a). Furthermore, when trained at a resolution of 256256, (b′) is not only comparable to (a) but also significantly faster.
@@
1) token representations with large receptive fields and reduced spatial redundancy can be utilized at the early stages. 2) It can diminish the feature map size by up to 16 times, leading to a significant reduction in memory access costs. 3) due to the aggressive stride design, there’s only a mild decrease in throughput when the resolution is increased, leading to effective performance enhancement (as shown in Fig. 2(b′), 8, and Tab. 2).
@@ 2.2 Analysis of Redundancy in Micro Design
First of all, in the early stages (Fig. 3 (a)), we observe that the top-performing heads tend to operate in a convolution-like manner, while heads that have minimal impact on performance when removed typically process information more globally. Also, as shown in Fig. 2 (b′′), the model using attention layers in the first stage exhibits a less favorable speed-accuracy trade-off compared to those employing depthwise convolution layers in the first stage. Hence, for efficiency, we use convolutions with spatial inductive bias as the token mixer in the initial stage.
@@ 2.3 Single-Head Self-Attention
| (a) : Bottleneck | 10.6 | 802 | 27195 | 73.6 |
| (b) : Full channel | 10.4 | 724 | 26432 | 75.1 |
| (c) : Partial channel (Ours) | 11.4 | 951 | 26878 | 75.2 |
@@ 3.1 Implementation Details
We conduct image classification on ImageNet-1K [23], which includes 1.28M training and 50K validation images for 1000 categories. All models are trained from scratch using AdamW [57] optimizer for 300 epochs with a learning rate of 10-3 and a total batch size of 2048. We use cosine learning rate scheduler [58] with linear warmup for 5 epochs. Weight decays are set to 0.025/0.032/0.035/0.03 for SHViT-S1 to S4. For fair comparison, we follow the same data augmentation proposed in [31], including Mixup [59], random erasing [60], and auto-augmentation [61]. For 3842 and 5122 resolution, we finetune the model for 30 epochs with weight decay of 10-8 and learning rate of 0.004. Additionally, we assess throughput performance across various hardware platforms. We measure GPU throughput on an Nvidia A100 with batch size of 256. For CPU and , we evaluate the runtime on an Intel(R) Xeon(R) Gold 5218R CPU @ 2.10GHz processor , with batch size of 16 (using a single thread). For , we convert the models to ONNX [62] runtime format. Mobile latency is measured using iPhone 12 with iOS version 16.5. We export the models (batch size is set to 1) using CoreML tools [63] and report the median latency over 1,000 runs. We also validate our model as an efficient vision backbone for object detection and instance segmentation on COCO [25] with RetinaNet [64] and Mask R-CNN [26], respectively. All models are trained under 1 schedule (12 epochs) following [40] on mmdetection library [65].
@@ 3.3 SHViT on downstream tasks
As shown in the above results, the large-stride patchify stem with 3-stage reduces not only computational costs but also generates meaningful token representations, especially at higher resolutions. Furthermore, the marked performance gap with EfficientViT [27], using a similar macro design, proves the efficacy of our micro design choices.
@@ 4 Related Work
Also, recent works [34, 35, 36, 37, 38, 39, 75, 27, 76, 77] have demonstrated that numerous heads function in similar ways and can be pruned without notably affecting performance. EfficientViT [27] proposes feeding attention heads with different splits of the full channel to improve attention diversity. In addition, [76] presents a regularization loss for multi-head similarity, while [78] explores head similarity across different layers. As opposed to reducing multi-head redundancy, we design module with single-head configuration, which not only inherently prevents multi-head redundancy but also saves computation costs.
@@ 5 Conclusion
In this work, we have investigated redundancies at both the spatial and channel dimensions of the architectural design commonly used by many established models. We then proposed 1616 patch embeddings with 3-scale hierarchical representations and Single-Head Self-Attention to address the computational redundancies. We further present our versatile SHViT, built upon our proposed macro/micro designs, that achieves ultra-fast inference speed and high performance on diverse devices and vision tasks.
@@
Discussion. While our macro design is effective, there is a need for fine-grained (high-resolution) features to enhance performance further or to recognize small objects. Therefore, our future work focuses on devising cost-effective ways to utilize them. Integrating the single-head design into existing sophisticated attention methods will be an intriguing way to explore.
6 Acknowledgement
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (NO.RS-2022-00166109) and (2022M3J6A1084845).
Appendix A Comparison with Tiny Variants of Large-scale Models
We compare our model against tiny variants of established models in Tab. 7. Our model, when applied to higher resolutions, outperforms state-of-the-art models in terms of parameter and throughput. Compared to Swin-T [40], our is 0.3% inferior in accuracy but is 2.3 / 9.5 faster on the A100 GPU / Intel CPU.
In Fig. 9, we also provide further results of Section 3.2. It demonstrates improved speed performance when increasing the resolution not only on mobile devices but also on other inference platforms compared to the recent models [8, 7]. This result showcases that our model can be a competitive alternative in real-world applications. Further analysis of these performance enhancements will be detailed in the following section.
Appendix B Memory Efficiency Analysis
Our model has a larger number of parameters compared to lightweight models. For instance, SHViT-S3 has 2.7 more parameters compared to EfficientNet-B0 [24]. However, an important consideration for deploying the model on resource-constrained devices is the memory access cost of the feature maps. On an I/O bound devices, the number of memory access for a given layer is as follows
Particularly, when increasing batch size to enhance throughput, or for applications that require high-resolution input, the impact of the first term in the above equation becomes significantly more critical. Our proposed macro and micro designs considerably reduce memory usage by eliminating redundancies in the first term’s and components, respectively. In Tab. 8, our model, despite having more parameters than EfficientNet-B0, consumes less test memory. Notably, the disparity in memory usage grows with increasing batch sizes.
Appendix C Further Results on COCO Detection
We also present results on COCO object detection benchmark [79] with DEtection TRansformer (DETR) [80, 81] framework in Tab. 9.
| Model | Params | FLOPs | Throughput (image/s) | Top-1 | |
| (M) | (G) | GPU | (%) | ||
| CaiT-XXS36 [82] | 17 | 3.8 | 1394 | 24 | 79.1 |
| Twins-PCPVT-S [49] | 24 | 3.8 | 3800 | 53 | 81.2 |
| Swin-T [40] | 28 | 4.5 | 2868 | 33 | 81.3 |
| TNT-S [83] | 24 | 5.2 | 1554 | 37 | 81.5 |
| CoAtNet-0 [84] | 25 | 4.2 | 2448 | 53 | 81.6 |
| DeiT-B [31] | 87 | 17.6 | 3227 | 21 | 81.8 |
| XCiT-S12 [85] | 26 | 4.8 | 3110 | - | 82.0 |
| PVTv2-B2 [70] | 25 | 4.0 | 2924 | 14 | 82.0 |
| FocalNet-T [86] | 28 | 4.4 | 2808 | 68 | 82.1 |
| ConvNeXt-T [87] | 29 | 4.5 | 3325 | 49 | 82.1 |
| SHViT-S4 | 17 | 1.0 | 14283 | 509 | 79.4 |
| 17 | 2.2 | 6702 | 315 | 81.0 | |
| 17 | 4.0 | 3957 | 198 | 82.0 | |
| Model | Top-1 | Params | Inference Memory (MB) / Throughput (images/s) | |||
| (%) | (M) | bs1 | bs32 | bs256 | bs1024 | |
| SHViT-S3 | 77.4 | 14.2 | 1855 / 147 | 1963 / 4691 | 2613 / 20522 | 5525 / 22309 |
| EfficientNet-B0 | 77.1 | 5.3 | 1931 / 175 | 2015 / 5427 | 3861 / 8433 | 10493 / 8706 |
The encoder of DETR consists of self-attention and FFN, and the decoder consists of self-attention, cross-attention, and FFN. To demonstrate the efficacy of our single-head attention module not only as a feature extractor but also as a detection head, we apply single-head design to the encdoer’s self-attention and decoder’s cross-attention layers. These two layers involve significant computational costs, thus employing a single-head design can greatly enhance the model speed. However, in the detection head, each of the attention weights localizes different extremities [88], making it challenging to simply combine them into a single-head design. Furthermore, we find that the multi-head design in both the initial layer and latter layers of the encoder/decoder is vital. Thus, we employ single-head attention modules in the 2nd, 3rd, and 4th layers of each encoder/decoder. To minimize performance degradation, we also increase the head dimension in the single-head module from 32 to 64. We train our model using the training recipe of Deformable DETR [81, 80]. As shown in Tab. 9, single-head module demonstrates reasonable performance as a detector head and is a competitive alternative for applications where inference speed is crucial.
| Method | Params | FPS | AP | AP50 | AP75 | APS | APM | APL |
| Deformable DETR w/ single-head | 37.1M | 31.4 (24% ) | 43.1 | 62.7 | 46.6 | 26.3 | 46.6 | 57.2 |
| Deformable DETR | 40.0M | 25.4 | 43.8 | 62.6 | 47.7 | 26.4 | 47.1 | 58.0 |
Appendix D More Details on Redundancy Experiments
In this section, we provide implementation details of section 2.2.
head similarity analysis. For each layer , the average cosine similarity value is computed as:
where is the number of heads. Then, the value is averaged for all batches.
head ablation study. In order to perform head ablation experiments, we modify the formula for Multi-Head Self-Attention (MHSA):
where the are mask vaiables with values in {0, 1}. When all are equal to 1, the above layer is equivalent to the MHSA layer. In order to ablate head , we simply set . We conduct experiments by selectively removing one or more attention heads from a given architecture during test time and assessing the resulting impact on accuracy. And we report the best accuracy for each layer in the model, i.e. the accuracy achieved by reducing the entire layer to the single most important head.
We further investigate head redundancy in DeiT-S-Distill [31], a vision transformer distilled with knowledge from ConvNets. In the distilled model, we can also observe a significant computational redundancy among many heads in the latter stages. Additionally, in the early stages, where many heads operate similarly to convolution, there is a relatively substantial decline in performance.
Appendix E Further Discussions on Related Works
About Macro Design. Our patch embedding scheme is similar to that of [10, 27], but the derivation process takes place from a completely different perspective. While [10] indirectly determines the patch embedding size through experiment grafting ResNet and DeiT, our work, on the other hand, investigate redundancy from the beginning, analyzing it separately in terms of spatial and channel. This allows us to address not only the spatial redundancy in traditional patch embedding but also propose a SHSA module, in contrast to [10] which employs MHSA (at mobile, SHViT-S4 80.2%/1.6ms vs. LeViT-192 80.0%/28.0ms). To the best of our knowledge, none of existing works have analyzed the effects (speed, memory efficiency) of resolving spatial redundancy in diverse environments (devices, tasks).
About Partial Design in SHSA. Partial channel design has also been employed in previous research [66, 30]. However, our work is distinct in both motivation and effectiveness. While prior work primarily focused on FLOPs (or throughput) and so employs convolutions (either depthwise or vanilla) on partial channels, this paper addresses multi-head redundancy by employing attention with single-head on partial channels. Furthermore, our SHSA, with preceding convolution, memory-efficiently leverages two complementary features in parallel within a single token mixer [12, 43].